Examine if the following eqns have any solution or not.
1) |x| = x+5
2) |x| = x -5

Answer by stanbon(75887)

1) |x| = x+5
For x>=0, x=x+5 has no solution
for x<0, -x=x+5, -2x=5, x= -5/2 is a solution.
----------
2) |x| = x -5
For x>=0, x=x-5 has no solution
For x<0, -x=x-5, -2x=-5, x=5/2
But that contradicts x<0; So, no solutions.
------------
==========
Cheers,
Stan H.